Fermi National Laboratory
 

Accelerator Update for February 19, 2007 - February 21, 2007

 

Monday February 19

The day shift began with Operations monitoring the Recycler stash (337.2E10), with the Antiproton Source (Pbar) stacking (12.7E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

Operations put the TeV into shot setup at 8:18 AM.

At 9:21 AM, Operations reported that cooling for shot setup was slow. An operator contacted an expert and then adjusted the beam current for the Electron Cooling.

At 9:30 AM, a Booster RF station (BRF6) tripped off. It reset.

At 9:33 AM, Operations reloaded protons into the TeV because the Recycler antiprotons were having emittance problems. An operator readjusted the beam current for Electron Cooling.

Operations put shot setup on hold at 9:46 AM due to more Recycler emittance problems.

TeV experts began a 36x0 Tevatron Electron Lens (TEL2) study at 10:17 AM.

The TeV study ended at 10:49 AM. Operations put the TeV through a dry squeeze.

The Neutron Therapy Facility (NTF) reported at 11:12 AM that they had completed treating their last patient for the day.

Operations put the TeV into shot setup at 12:09 AM.

Operations turned off the Booster at 1:01 PM.

Operations reported at 1:14 PM that the Pbar AP1 enclosure was ready for access.

Operations began preparing Booster for access at 1:27 PM.

Operations established store 5236 at 1:43 PM with an initial luminosity of 173.15E30.

Operations began preparing MI for access at 2:06 PM.

MI came out of access at 4:28 PM.

Booster had come out of access and Operations began turn on procedures for MI and Booster at 4:31 PM.

At 5:30 PM, Operations reported that BRF16 had another problem that required a tunnel access. Booster RF experts went into access, fixed BRF16, and came out at 6:05 PM.

Operations had established beam to NuMI and Pbar by 6:33 PM.

At 6:44 PM, Operations reported that BRF14 was down and needed a tunnel access to repair. Operators bypassed the station.

Operations reported at 11:56 PM that the Recycler had stashed antiprotons twice during the evening shift.

Tuesday February 20

The midnight shift began with Operations monitoring store 5236 (58.25E30) and the stash (108.7E10), with Pbar stacking (20.1E10), and with NuMI taking beam.

At 1:50 AM, BRF6 tripped off for the first of six times in five hours. It reset each time.

At 4:41 AM, Operations reported that a transfer of antiprotons from Pbar didn’t make it into MI due to a problem with an Accumulator kicker (A:EKIK).

At 5:50 AM, another transfer from Pbar was lost, and at 7:11 AM yet another was lost due to the kicker. Operations contacted a Pbar expert who said that the kicker was firing one revolution early. A power supply expert investigated.

A Meson/MTest expert established beam to MTest at 7:16 AM.

At 7:59 AM, the NuMI beam permit tripped off due to a bad beam position on their target. It reset.

Operations reported at 8:37 AM that the Recycler had stashed three times during the midnight shift.

At 9:30 AM, Operations reported that the Linac Buncher had gone out of phase for the first of three times in 33 minutes. A Linac expert responded and reset the device each time.

A TeV expert began a TEL1 study at 10:01 AM.

A TeV expert began a crystal collimator study at 11:42 AM.

A MI expert began a slipstacking study at 1:26 PM.

TeV experts completed the crystal collimator study at 2:07 PM and then began a Schottky study.

TeV experts completed their studies. Operations terminated store 5236 at 2:31 PM.

Operations prepared CDF and D0 for access at 2:44 PM.

Operations reported at 4:10 PM that the Recycler had stashed three times during the day shift.

CDF and D0 were both out of access by 5:20 PM.

Operations put the TeV into a dry squeeze at 5:37 PM.

At 5:44 PM, LRF5 tripped off. Operators brought the station up slowly and beam resumed at 5:59 PM.

Operations put the TeV into shot setup at 6:59 PM.

Operations established store 5238 at 8:52 PM with an initial luminosity of 243.91E30.

Operations reported at 11:52 PM that the Recycler had stashed twice during the evening shift.

Wednesday February 21

The midnight shift began with Operations monitoring store 5238 (141.14E30) and the stash (61.8E10), with Pbar stacking (54.4E10), and with NuMI taking beam.

At 12:53 AM, LRF5 tripped off for the first of three times in 25 minutes. It reset.

At 1:44 AM, LRF5 tripped off again. Operators had it back up within eight minutes.

At 2:52 AM, LRF5 tripped off. It took operators five minutes to bring the station back on this time.

At 2:59 AM, LRF5 tripped off for the first of three times in 10 minutes. Eventually it reset and ran.

At 3:22 AM, BRF6 tripped off. It reset.

At 4:18 AM, Operations reported that the H- Source had been sparking.

At 6:03 AM, KRF6 tripped off. It reset

At 7:12 AM, BRF15 & 16 tripped off due to a vacuum high-pressure indication. Operators investigated and found no problems. They reset the stations.

Operations reported at 8 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Wednesday

The plans for today are to stack and stash. At 8:46 AM, Fermilab suffered from a ComEd power glitch that resulted in losing store 5238, a TeV quench at sector C1, and many devices tripping off. Fortunately, the stash and stack remained intact. We intend to recover everything and establish a store by this afternoon.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/21/2007    email Fermilab